OpenStudy (anonymous):

What is the point slope form of an equation with the slope 3/5 that passes through the point (10,2)?

OpenStudy (amistre64):

point slope form: \[y-2 = \frac{3}{5}(x-10)\]

OpenStudy (amistre64):

\[y-Py = m(x-Px)\]

OpenStudy (anonymous):

yah you are correct. mistaken

OpenStudy (anonymous):

amistre got your signs right? you sure its + or -?

OpenStudy (anonymous):

The signs are correct. To see this consider what the slope of a line means. Slope of a line is rise/run or \[m = \Delta y / \Delta x\] Or m= y-2/x-10 where (2,10) is a given point on the line and x, y is any other point on the line. We know the slope is 3/5 so our equation become 3/5=y-2/x-10 which rearranges to point-slope form as y-2=3/5(x-10)

OpenStudy (anonymous):

well i have a booklet and the answer he gave is incorrect, there are 2 similar options and the signs are mixed, the + and -

OpenStudy (anonymous):

If the given point is (+2,+10) and slope is m=+3/5 then the answer we gave is correct (and your booklet is wrong). Are you sure of the signs on the given point?

OpenStudy (anonymous):

ooops given point should read (+10, +2) above